1. The selling price of a television is $ 1100 and the cost to the retailer is $225. What is the​ retailer's gross profit from the sale of the​ television?

A) $1100

B) $875

C) $0

D) $225

2. A company has a beginning inventory of $ 60,000 and purchases during the year of $ 120,000. The beginning inventory consisted of 2,000 units and 9,000 units were purchased during the year. The company has 5,000 units left at year end. -Under average ​cost- , what is Cost of Goods​ Sold? (Round any intermediary calculations to two decimal places and your final answer to the nearest​ dollar.)

120,000

98,160

180,000

147,240

3. Tomasino's inventory records show the following data at January​ 31:

Beginning inventory Jan. 1            120 units at $ 5 per unit

Jan. 10 purchase                              330 units at $ 12 per unit

Jan. 22 purchase                              110 units at $ 13 per unit

At January​ 31, 200 units are still on hand. What is the cost of the ending inventory at January 31 if Tomasino uses the FIFO​ method?

  1. 1560
  2. 2510
  3. 1000
  4. 1770

4. ​Thelen's inventory records show the following data at January​ 31:

Beginning inventory Jan. 1                        90 units at $ 6 per unit

Jan. 10 purchase                                         320 units at $ 11 per unit

Jan. 22 purchase                                         120 units at $ 12 per unit

At January​ 31, 200 units are still on hand. What is the cost of the ending inventory at January 31 if Thelen uses the LIFO​ method?

A) 1760

B) 1200

C) 2400

D) 1860
Answer #1

Gross Profit = Sales – cost of goods sold

= 1100-225

= $875

i.e. B

Cost of goods sold = Average cost per unit*Units sold

Average cost = Total cost/Total units

= (60,000+120,000)/11,000

= $16.36

Cost of goods sold = 16.36*6,000

= $98,160

FIFO method states that goods purchased first will be sold first

Ending inventory = 110*13 + 90*12

= $2510

i.e. B

LIFO method states that goods purchased later will be sold first

Ending inventory = 90*6 + 110*11

= $1750